OpenStudy (anonymous):

Will Medal and Fan! Please help. I have the answer. Find the complete solution (to the nearest tenth) if sin𝛳= -.98544973 and 0≤𝛳≤2π.

OpenStudy (anonymous):

@mathmale please help.

OpenStudy (welshfella):

first use your calculator to find the angles whose sine is 0.98544973 ( set angles to radians)

OpenStudy (welshfella):

can you do that? the required key will be either sin-1 or arcsin

OpenStudy (welshfella):

r u there?

OpenStudy (anonymous):

yes i was trying to find free TI 84 calculator online to use

OpenStudy (anonymous):

so press sin-1(0.98544973 ?

OpenStudy (anonymous):

I got 1.4

OpenStudy (anonymous):

is that right? @welshfella

OpenStudy (anonymous):

thanks I got it now @welshfella it's 1.4+π =4.5 -1.4+2π=4.9